OpenStudy (anonymous):

Prove that the following identity is true. (1/1 + cos x) + (1/1 − cos x) = 2 csc^2

OpenStudy (anonymous):

add

OpenStudy (anonymous):

\[\frac{1}{1+a}+\frac{1}{1-a}=\frac{1+a+1-a}{1-a^2}\]

OpenStudy (anonymous):

Ok What confuses me is the part I got incorrect in the image.

OpenStudy (anonymous):

ooh because you made a mistake in the third line

OpenStudy (anonymous):

Yea, I have no idea how to get that part.

OpenStudy (anonymous):

the machine thinks you are an idiot, so instead of giving you the whole denominator to fill in, it already filled in part of it for you

OpenStudy (anonymous):

on no, you are right. damn

OpenStudy (anonymous):

lol, hmm... that weird. Ok Thanks.

OpenStudy (anonymous):

i mean \(1-\sin^2(x)\)

OpenStudy (anonymous):

Hmm.. it didn't work. That was my last shot at it lol.

OpenStudy (anonymous):

Thanks though